OpenStudy (anonymous):

Your piggy bank has a total of 40 coins in it; some are dimes and some are nickels. If you have a total of $3.25, how many dimes do you have? 30 25 20 15

OpenStudy (anonymous):

set up two equations. one for the coins 40= x+y one for the amount 3.25 = .05 + .10y

OpenStudy (anonymous):

to make things easier for the amount multiply through by 10 just to get rid of the annoying decimals. it wont mess anything up.

OpenStudy (anonymous):

it's a system of equations 40 = d + n 3.25 = .05n + .10d

OpenStudy (anonymous):

25

OpenStudy (anonymous):

solution: 25,15 Dimes, Nickels

OpenStudy (anonymous):

I got 15